OpenStudy (imsleeziboii):

Will give medal and fan 2 questions.

OpenStudy (thadds2003):

With.....?

OpenStudy (imsleeziboii):

To simplify the radical below, which of the following expressions would be multiplied by the radical? \[\frac{ -2\sqrt{3} }{ 3\sqrt{2} }\]

OpenStudy (thadds2003):

Ok, Give me one sec.....

OpenStudy (imsleeziboii):

oky

OpenStudy (imsleeziboii):

Do you want the choices or no?

OpenStudy (thadds2003):

No, I got dis....

OpenStudy (imsleeziboii):

Okay lmao

OpenStudy (thadds2003):

Wich subject is dis?

OpenStudy (imsleeziboii):

Alegebra 1

OpenStudy (thadds2003):

Ok, so the first one is -0.816497. The second question, please.

OpenStudy (imsleeziboii):

Thats not one of the choices man, sorry lol.

OpenStudy (thadds2003):

Ok, well then give me the choices.

OpenStudy (thadds2003):

Shoot, sorry, cant do dis. Heres a medal tho. Hope some one helps, @mathmale @just_one_last_goodbye

OpenStudy (imsleeziboii):

Thanks for the attempt

OpenStudy (imsleeziboii):

@mathmale

OpenStudy (thadds2003):

hey, bump your question to the top of the list.

OpenStudy (imsleeziboii):

okay I did

OpenStudy (imsleeziboii):

@mathstudent55

OpenStudy (thadds2003):

Wait, I figured out the answer.

OpenStudy (imsleeziboii):

?

OpenStudy (thadds2003):

\[\frac{1}{\sqrt[3]{2}\cdot \sqrt{3}}\]

OpenStudy (imsleeziboii):

That isnt one of the choices either lmao

OpenStudy (imsleeziboii):

but thanks for trying man

OpenStudy (imsleeziboii):

i tried that too and it didnt find the right one

OpenStudy (thadds2003):

Oh, sorry.

OpenStudy (imsleeziboii):

Its all good bro